body of mass 7 kg moves in a (counterclockwise) circular path of radius 4 meters, making one revolution every 13 seconds. You may assume the circle is in the xy-plane, and so you may ignore the third component.

how do you find the centripetal force acting on the body

how do you find the magnitude of the force??
Answer #1
centripetal acceleration = Ac = v^2/R = w^2 R
where w (omega) is the angular velocity in radians/second

here w = 2 pi radians/13 seconds
so
w^2 R = (4 pi^2/169) 4
then F = m a = 7 (4 pi^2/169) 4
answered by: Violentine